A rectangle has a perimeter (P) of 58 inches. The length (L) is one more than 3 times the width (W). write and solve a system of linear equations to find the length and width of the rectangle.
Answer by blwinbbbles(106)

First write equations for what you know: A perimeter is the distance around the rectangle, there are 2 lengths and 2 widths, we know the perimeter is 58 inches..So...
2L + 2W = 58 this is equation 1
Next we know that the length is one more than 3 times the width..so..
L = 3W + 1 equation 2
Since we know that L = 3W + 1 we can substitute the L in equation 1
2(3W + 1) + 2W = 58 solve the problem
6W + 2 + 2W = 58 add the common terms
8W + 2 = 58 subtract 2 from both sides
8W = 56 divide by 8
W = 7
Now that you know W, go back and solve for L....use equation 2
L = 3(7) + 1
L + 21 + 1
L = 22
hope this helps